OpenStudy (anonymous):

I know that xy+3x+2y+6 factorised equals (x+2)(Y+3), but i'm not sure how to get that?

hero (hero):

Let's see.... xy + 3x + 2y + 6 = x(y + 3) + 2(y + 3) = (y + 3)(x + 2) = (x + 2)(y + 3)

OpenStudy (anonymous):

Thanks!

hero (hero):

yw